Find the roots of the following quadratic equations by factorisation: 
100x ^2 – 20x + 1 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, 100x^2 – 20x + 1=0
Taking LHS, = 100x^2 – 10x - 10x + 1
= 10x(10x - 1) -1(10x - 1)
= (10x - 1)^2
The roots of this equation,
100x^2 – 20x + 1=0,
are the values of x for which (10x - 1)^2= 0
∴ (10x - 1) = 0
or (10x - 1) = 0
⇒x = 1/10 or x = 1/10

Thanks